Can someone help me solve this step by step using dimensional analysis. And explain how please. Thank you

8. Amoxil 500ms po o6 h is ordered for a child weighing 50 kg The recommended high dosage of Amoxil is 80 mg to 100 mglk/24 h q12. a. What is the single dose range for this child? b. Is the order doe safe to administer? Why?
8. Amoxil 500mg po q h is ordered for a chid weighing 50kg. The recommended high dosage of Amoxil is 80 mg to 100 mglkg24 h q12. a. What is the single dose range for this child? b. Is the order doe safe to administer? Why?
Answer #1

a) This child can be given 4000-5000mg/24hrs

2000-2500mg can be given 12hrly

Calculation- Recommended High dosage of Amoxil is 80-100mg

childs weight is 50kg

so 50×80=4000 and 50×100=5000

Maximum single dose range is 4000-5000mg

b) Yes the does ordered is safe to administer as the maximum dose is 4000-5000mg/24h (as shown in above calculation) and the dose ordered is 2000mg/24h (500/6hrly so 500×4=2000mg/24h) which is much less than the recommended high dose of amoxil.